OpenStudy (trumpdonald):

PLEASE ITS AN EMERGENCY MY MOM WANTS TO KILL ME What is the equation, in point-slope form, for a line that goes through (8, −4) and has a slope of −56 ? y−4=−56(x+8) y+4=−56(x−8) y+4=−56(x+8) y−4=−56(x−8)

OpenStudy (trumpdonald):

HELPPPPPPPP

OpenStudy (s4sensitiveandshy):

come on trump

OpenStudy (trumpdonald):

? HELP PLZ

OpenStudy (s4sensitiveandshy):

ok mr. Trump :)

OpenStudy (s4sensitiveandshy):

\[y-y_1=m(x-x_1)\] point slope form \[(x_1,y_1)\] is the point and m is slope.

OpenStudy (trumpdonald):

SO THATS THE ANSWER

OpenStudy (s4sensitiveandshy):

No given information slope = -56 the point (x_1,y_1) = ( 8 ,-4) plug these into the equation

OpenStudy (trumpdonald):

give me the answer now

OpenStudy (trumpdonald):

plz

OpenStudy (s4sensitiveandshy):

don't give me orders

OpenStudy (trumpdonald):

sorry

OpenStudy (trumpdonald):

my mom wants to kill me really

OpenStudy (s4sensitiveandshy):

all you need to do is replace *m* with -56 x_1 with 8 and y_1 with -4 that's it :)

OpenStudy (trumpdonald):

she has a knife

OpenStudy (trumpdonald):

watching me

OpenStudy (trumpdonald):

if i dont get the answer in 30 sec she will kill me

OpenStudy (s4sensitiveandshy):

counting

OpenStudy (trumpdonald):

nooooooooooooooooooooooo plzzzzzzzzz

OpenStudy (trumpdonald):

oh noooo she is killing meeeeeeee

OpenStudy (s4sensitiveandshy):

you are wasting your time. do you know how to replace one variable with a number ?

OpenStudy (trumpdonald):

no im not smert

OpenStudy (s4sensitiveandshy):

yes you are.

OpenStudy (trumpdonald):

ok ill try just use stupid people words

OpenStudy (trumpdonald):

@S4Sensitiveandshy

OpenStudy (trumpdonald):

@S4Sensitiveandshy

OpenStudy (trumpdonald):

@S4Sensitiveandshy

OpenStudy (trumpdonald):

@3mar

OpenStudy (s4sensitiveandshy):

\[y-y_1=m(x-x_1)\] this is how we should replace *m* with *-56* \[y-y_1=-56(x-x_1)\] do the same thing with point (8,-4)

OpenStudy (3mar):

I can't help that Racist killer @TrumpDonald!

OpenStudy (trumpdonald):

y-ysmall1 = -56(8,-4)

OpenStudy (trumpdonald):

it just a name haha

OpenStudy (3mar):

name represents that sick man!

OpenStudy (s4sensitiveandshy):

\[y-y_1 =-56(x-x_1)\] (8,-4) represent the point (x_1,y_1) so replace x_1 with 8 and y_1 with -4

OpenStudy (trumpdonald):

im sorry its just a joke

OpenStudy (s4sensitiveandshy):

(x,y) point where first number is x-coordinate and the 2nd one is y-coordinate

OpenStudy (trumpdonald):

assalam u alykum im a muslim

OpenStudy (trumpdonald):

i am :/

OpenStudy (trumpdonald):

oh so is b right?

OpenStudy (s4sensitiveandshy):

ok nice so what would you get ?

OpenStudy (trumpdonald):

b?

OpenStudy (s4sensitiveandshy):

yes that's right.

OpenStudy (3mar):

Muslim and put that rubbish to represent you!!!!!!! It does not make any sense!

OpenStudy (trumpdonald):

trump is a meme he is a really bad president and i do not vote for him, its all a joke please dont be offened

OpenStudy (trumpdonald):

wallah ani muslima isme saara

OpenStudy (3mar):

so why this photo is for you? Do you accept that you are represented by this Hater of women?

OpenStudy (trumpdonald):

its a joke :) i do not like trump but i dont hate him but you gotta say he is SUCH a joke

OpenStudy (s4sensitiveandshy):

Khalas.

OpenStudy (trumpdonald):

zalagna

OpenStudy (trumpdonald):

wow how did 3 muslims get in one question? i guess theres more of us then i thought

OpenStudy (s4sensitiveandshy):

3?

OpenStudy (trumpdonald):

me you and 3mar

OpenStudy (trumpdonald):

oh your not a muslim?

OpenStudy (s4sensitiveandshy):

I know few words in Arabic.

OpenStudy (trumpdonald):

ooooohhhhh